OpenStudy (anonymous):

your uncle is three years older than your aunt. the product of their ages is 3190. find the ages of your aunt and uncle.

OpenStudy (anonymous):

Well, let the aunt's age be x. Then the uncle's age is x+3, and we have that \[x(x+3) = 3190\] \[x^2 + 3x -3190 = 0\] \[x=\frac{-3 \pm \sqrt{9+4(3190)}}{2}\] \[x = \frac{-3 \pm 112.96017}{2}\] At this point note that the age of the aunt cannot be negative, so disregard the negative solution and keep the positive: \[x = 54.98008\] and this is the aunt's age. The uncle is 57.98008

OpenStudy (anonymous):

How old are my aunt and uncle if my uncle is 3 years older than my aunt and the product of their ages is 3190? ... Will my aunt and uncle get divorced because of me? ... If someone's aunt marries your uncle does this make you related?